use the slope intercept form to graph the equation x=4\5 y
Answer by ewatrrr(24785)

 
Hi,
x=4\5 y
y = (5/4)x, slope = (5/4) y changes 5, x has a corresponding change of 4
(0, 0)
(4, 5)
(-4,-5)